OpenStudy (rayford):

Rank from most conductive to least conductive. To rank items as equivalent, overlap them. POTASSIUM BROMINE SUGAR BARIUM CHLORIDE BUTANIOC ACID

OpenStudy (boldjon):

Well, I did least to most Ions are what conduct electricity in solution. Sugar does not ionize - lousy conductivity Benzoic is a weak acid, ie it does not ionize much KCl will dissociate totally into two ions per molecule AlCl3 will give more ions

OpenStudy (anonymous):

MgBr2(s) + ∞H2O → Mg^2+(aq) + 2Br^-(aq) three ions per mol dissolved hence 0.60 mol of ions produced C6H12O6(s) + ∞H2O → C6H12O6(aq) dissolves by H-bonding interactions and remains undissociated 0 ions produced Let A = C6H5COO HA + H2O ⇋ H3O^+ + A^- pKa [1] = 4.202 = 6.28×10^-5 put into Ka Ka = [H3O^+][A^-]/[HA] = x^2/0.2 = 6.28×10^-5 x = 3.5×10^-3 so #of ions is twice this 7×10^-3 KBr(s) + ∞H2O → K^+(aq) + Br^-(aq) two ions per mol dissolved hence 0.40 mol of ions produced. Conductivity MgBr2 > KBr > C6H5COOH > C6H12O6

OpenStudy (anonymous):

just google this one lol

OpenStudy (rayford):

thanks lol...